OpenStudy (brittany):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (curry):

y = 3x-3

OpenStudy (anonymous):

y = 3x + b 9 = 12 + b b = -3 y = 3x - 3

OpenStudy (anonymous):

y = 3x - 3

OpenStudy (curry):

leave medal

hero (hero):

pk posts full solutions, so he's the fastest

OpenStudy (curry):

lol

OpenStudy (anonymous):

I try to explain, if you want answers, go to wolfram alpha.

OpenStudy (curry):

i happen to have a slow compuiter